# FedMKT: Federated Mutual Knowledge Transfer for Large and Small Language Models

### v2

@@ Abstract
Recent research in federated large language models (LLMs) has primarily focused on enabling clients to fine-tune their locally deployed homogeneous LLMs collaboratively or on transferring knowledge from server-based LLMs to small language models (SLMs) at downstream clients. However, a significant gap remains in the simultaneous mutual enhancement of both the server’s LLM and clients’ SLMs. To bridge this gap, we propose FedMKT, a parameter-efficient federated mutual knowledge transfer framework for large and small language models. This framework is designed to adaptively transfer knowledge from the server’s LLM to clients’ SLMs while concurrently enriching the LLM with clients’ unique domain insights. We facilitate token alignment using minimum edit distance (MinED) and then selective mutual knowledge transfer between client-side SLMs and a server-side LLM, aiming to collectively enhance their performance. Through extensive experiments across three distinct scenarios, we evaluate the effectiveness of FedMKT using various public LLMs and SLMs on a range of NLP text generation tasks. Empirical results demonstrate that FedMKT simultaneously boosts the performance of both LLMs and SLMs.
@@
Tao Fan1, 2, Guoqiang Ma2, Yan Kang2, Hanlin Gu2, Yuanfeng Song2, Lixin Fan2, Kai Chen1, Qiang Yang1,2 1 Hong Kong University of Science and Technology, China 2 WeBank, China Correspondence: tfanac@cse.ust.hk
@@ 1 Introduction
Large Language Models (LLMs) have emerged as a transformative force in artificial intelligence, profoundly altering our perception of natural language processing capabilities. The advent of cutting-edge LLMs like ChatGPT  OpenAI (2022), and LLaMA Touvron et al. (2023) with their billions of parameters, has sparked the imagination of both researchers and practitioners, owing to their exceptional performance across diverse text generation tasks. Despite their widespread success in various general NLP tasks, LLMs face challenges that hinder their adoption in domain-specific applications Kang et al. (2023) Fan et al. (2023). The primary challenges include domain-specific knowledge Privacy, constrained computing resources, and mutual knowledge transfer between the LLM and SLMs. A significant challenge arises from the inherent model heterogeneity between the LLM and SLMs, particularly when aligning distributions of output logits. The mismatch between the tokenizers of different LLM and SLMs poses a notable obstacle. Furthermore, the mutual transfer of knowledge between the server’s LLM and clients’ SLMs remains a largely unexplored area in academic literature, warranting further investigation.
To fill these gaps, we propose FedMKT, a novel federated mutual knowledge transfer framework designed to enhance the performance of both large and small language models. By leveraging the complementary strengths of federated learning and knowledge distillation, FedMKT facilitates effective mutual knowledge transfer between clients’ SLMs and the LLM owned by the server.
As illustrated in Figure 1, FedMKT deploys an LLM on the server and a set of $K$ heterogeneous SLMs across various clients. The cornerstone of FedMKT lies in its selective mutual knowledge transfer process. During each round of federated learning, the clients transmit the output logits of their updated SLMs on the public dataset to the server. The server then selectively aggregates and distills the knowledge encoded within these SLMs output logits into the server-side LLM. This process allows the server LLM to incorporate the domain-specific knowledge learned by the clients, thereby enhancing its comprehensive capabilities. Simultaneously, the server-side LLM also selectively distills its knowledge to the clients’ SLMs, which is similar to the knowledge transfer from clients to the server. By leveraging the knowledge of the server LLM, the clients’ SLMs are able to improve their performance and generalize better to unseen data. To address the model heterogeneity between the LLM and SLMs, FedMKT incorporates a token alignment technique utilizing minimum edit distance (MinED) prior to knowledge transfer. This alignment ensures seamless integration and efficient knowledge transfer between LLM and SLMs.
@@
  • Federated Mutual Knowledge Transfer Framework. FedMKT introduces a novel federated mutual knowledge transfer framework that enables effective knowledge transfer between an LLM deployed on the server and SLMs residing on clients. This framework fills the gap by simultaneously enhancing both the server’s LLM and the clients’ SLMs.
@@
  • Empirical Evaluation and Performance Enhancement. Extensive experiments conducted based on various publicly available LLMs and SLMs demonstrate the competitive performance of FedMKT across a wide range of NLP text-generation tasks. We evaluate FedMKT with heterogeneous, Homogeneous, and One-to-One settings. The results show that the performance of SLMs can be significantly enhanced with the help of the LLM, while the LLM can deliver comparable results to fine-tuning with all clients’ data centralized.
@@ 2.1 Model Heterogeneous Federated Learning
Model heterogeneous federated learning (MHFL) aims to address the challenges associated with heterogeneity in federated learning. Initial research in MHFL primarily concentrated on addressing heterogeneity in model architectures. Various approaches have been proposed to accommodate clients with different model architectures participating in a federated learning task. These methods typically involve techniques such as knowledge distillation Hinton et al. (2015), mutual learning and split learning that can handle heterogeneous models. Knowledge distillation-based MHFL methods, such as FedMD Li and Wang (2019) and FedET Cho et al. (2022), involve the server aggregating the output logits of different clients’ heterogeneous models on a public dataset to construct global logits. Mutual learning-based MHFL, such as Deep Mutual Learning (DML) Zhang et al. (2018), PFML Yang et al. (2021) and FedLoRA Yi et al. (2023), design a small homogeneous model and a large heterogeneous model in each client. Split learning-based MHFL approaches, such as FedClassAvg Jang et al. (2022) and CHFL Liu et al. (2022), share a homogeneous classifier to improve model classification while personalizing the local feature extractor.
While previous works have mainly focused on computer vision scenarios, the literature has limitedly explored MHFL in LLMs. This gap motivates this study, which aims to explore MHFL in the context of LLMs.
@@ 2.2 Federated Learning for LLMs
Parameter-Efficient Fine-Tuning (PEFT) methods Houlsby et al. (2019); He et al. (2021); Lester et al. (2021); Li and Liang (2021); Hu et al. (2021) offer a direct solution to the issues of communication overhead and fine-tuning costs in federated learning (FL) for LLMs. A number of studies have built upon PEFT methods in the context of FL for LLMs, including FedPETuning Zhang et al. (2022b), Federated Adapter Tuning Cai et al. (2022), Federated Prompt Tuning Zhao et al. (2022), and FATE-LLM Fan et al. (2023). For example, the FedPETuning Zhang et al. (2022b) has demonstrated a significant reduction in communication overhead, reducing 1 to 2 orders of magnitude compared to full fine-tuning in the FL setting. These findings imply that FL clients, such as devices with limited storage capacity, can greatly benefit from PEFT methods. These methods enable the sharing of LLMs across different tasks while maintaining only a few parameters for each task, thereby reducing the storage requirement. By leveraging PEFT methods, FL clients can efficiently adapt LLMs to their specific needs while minimizing communication overhead and fine-tuning costs.
@@ 3 The Proposed FedMKT Method
In this section, we introduce FedMKT, an innovative and parameter-efficient federated mutual knowledge transfer approach for large and small language models. The FedMKT primarily comprises two key modules: Bidirectional Token Alignment and Selective Mutual Knowledge Transfer. We will elaborate on these two modules in Section 3.2 and Section 3.3, respectively after we define the problem we try to address in Section 3.1.
@@ 3.1 Problem Definition
We consider the federated learning setting, involving one server that owns an LLM $f_{\psi}$ parameterized by $\psi$ and $K$ clients that each client $k$ has an SLM $g_{\phi_{k}}$ parameterized by ${\phi}_{k}$. Each client owns a local private dataset $\mathcal{D}_{k}$ with $N$ training samples, and all clients and the server share a public dataset $\mathcal{D}_{p}$.
@@
FedMKT solves the optimization problems formulated in Eq.(1), Eq.(2), and Eq.(3) in an efficient and privacy-preserving manner. We illustrate the workflow of FedMKT in Figure 1 and elaborate on the associated training algorithm in Algorithm 1.
@@ 3.2 Bidirectional Token Alignment
To tackle this issue, we adopt dynamic programming techniques to promote robust alignment, as evidenced in studies Wan et al. (2024); Fu et al. (2023). Utilizing LLaMa2 and Bloom as illustrative examples, we establish an optimized vocabulary mapping table based on minimum edit distance (MinED). This mapping table identifies the closest Bloom token for each LLaMa2 token (e.g., ’utilize’ for ’util’). We then tokenize a sentence using both tokenizers and apply a dynamic programming algorithm to determine the optimal matching path. When multiple LLaMa2 tokens align to a single Bloom token (e.g., ’util’ and ’ize’ aligning to ’utilize’), we handle them according to the mapping table. Please refer to Appendix B for more details.
@@ 3.3 Selective Mutual Knowledge Transfer Between LLM and SLMs
where $\ell_{\text{CE}}$ is the cross-entropy loss; the model parameters $\psi$ and $\phi_{k}$ are frozen during fine-tuning.
Next, the server and clients conduct selective knowledge transfer to each other. The motivation for applying selective knowledge transfer is that some clients’ knowledge may adversely affect the performance of LLM on the server and vice versa in a heterogeneous environment. Therefore, it is critical to guarantee that the knowledge transferred between the server and clients is positive to the performance of LLM and SLMs. To this end, we propose a selective knowledge transfer strategy on both the server and client sides, termed DualMinCE.
@@
where $\lambda$ is the hyperparameter that controls the weight of mutual knowledge transfer.
@@ 4.1 Setup
Datasets. We evaluate FedMKT comprehensively on 6 QA datasets and 2 instruction-following datasets. Specifically, for QA tasks, we use RTE Wang et al. (2019), WTC Wang et al. (2019), BoolQ Clark et al. (2019), CommonsenseQA(CQA) Talmor et al. (2018), ARC-E Clark et al. (2018), ARC-C Clark et al. (2018) to evaluate FedMKT. As for instruction-following tasks, we evaluate FedMKT on S-NI Wang et al. (2022), DialogSum Chen et al. (2021).
@@
  • Centralized, in which the server’s LLM is fine-tuned locally using the datasets combining private datasets of involved clients and the public dataset. In the One-to-One setting, the data of one client and the public data are used to fine-tune the server’s LLM, whereas in other settings, the data of all four clients and the public data are used to fine-tune the LLM;
@@
  • Standalone, in which each client independently fine-tunes its local SLM using its private dataset;
  • FedAvg, representing the standard federated averaging algorithm. FedAvg is only used in homogeneous settings because it requires all clients’ models have the same architecture.
@@
Evaluation Metrics. For the QA datasets, we primarily use Accuracy as the evaluation metric, whereas for the instruction-following datasets, we primarily rely on Rouge-L.
@@ 4.2 Evaluation on Heterogeneous Setting
Tables 2 show that FedMKT performs superior over Zero-Shot and Standalone on all clients’ SLMs. Take the RTE dataset as an example, FedMKT outperforms Zero-Shot by 34% and Standalone by 7% on the GPT-2-xlarge SLM; FedMKT surpasses Zero-Shot by 25% and Standalone by 5% on the OPT-1.3B SLM; FedMKT-SLM achieves a 17% improvement over Zero-Shot and a 6% improvement over Standalone on the Bloom-1.1B SLM; FedMKT-SLM outperforms Zero-Shot by 18% and Standalone by 6% on the LLaMa2-1.3B SLM. These empirical results demonstrate that, by leveraging FedMKT, SLMs are able to effectively leverage the knowledge transferred from the LLM, leading to enhanced model capabilities.
@@ 4.3 Evaluation on Homogeneous Setting
The bottom sub-table of Table 3 shows that FedMTK performs better than the Zero-Shot, Standalone, and FedAvg due to the assistance of the server’s LLM. For example, in the CQA dataset, FedMKT outperforms FedAvg by 4% on the LLaMa2-1.3 SLM and by 5% on the OPT-1.3B SLM, respectively.
@@ 5 Conclusions
In this study, we have presented FedMKT, a parameter-efficient federated mutual knowledge transfer framework tailored for large and small language models. FedMKT bridges the gap between the server-side LLM and clients’ SLM, enabling selective mutual knowledge transfer while preserving data privacy. Through extensive experiments across three distinct scenarios, we have demonstrated that FedMKT simultaneously boosts the performance of both LLMs and SLMs.
@@
In this study, we transfer knowledge between the server and clients using logits of a public dataset, motivated by efficiency and privacy considerations. Although empirical evidence suggests that sharing logits of public datasets between the server and clients is more privacy-preserving than sharing model gradients or parameters Li and Wang (2019); Cho et al. (2022), there is no theoretical guarantee that this approach does not compromise the privacy of clients’ sensitive data. This issue warrants further investigation. Furthermore, our study is limited by computational and storage constraints, which preclude the exploration of larger language models. This highlights the inherent trade-off between utility and efficiency. Our future research aims to investigate and optimize this trade-off.

### v4

@@ Abstract
Recent research in federated large language models (LLMs) has primarily focused on enabling clients to fine-tune their locally deployed homogeneous LLMs collaboratively or on transferring knowledge from server-based LLMs to small language models (SLMs) at downstream clients. However, a significant gap remains in the simultaneous mutual enhancement of both the server’s LLM and clients’ SLMs. To bridge this gap, we propose FedMKT, a parameter-efficient federated mutual knowledge transfer framework for large and small language models. This framework is designed to adaptively transfer knowledge from the server’s LLM to clients’ SLMs while concurrently enhancing the LLM with clients’ unique domain insights. We facilitate token alignment using minimum edit distance (MinED) and then selective mutual knowledge transfer between client-side SLMs and a server-side LLM, aiming to collectively enhance their performance. Through extensive experiments across three distinct scenarios, we evaluate the effectiveness of FedMKT by utilizing diverse public LLMs and SLMs on a variety of NLP text generation tasks. Empirical results demonstrate that FedMKT simultaneously boosts the performance of both LLMs and SLMs. Our code has been contributed to the FATE open-source project and is now publicly accessible at https://github.com/FederatedAI/FATE-LLM/tree/main/python/fate_llm/algo/fedmkt
@@
Tao Fan1, 2, Guoqiang Ma2, Yan Kang2, Hanlin Gu2, Yuanfeng Song2, Lixin Fan2, Kai Chen1, Qiang Yang1,2 1 The Hong Kong University of Science and Technology, Hong Kong, China 2 WeBank Co., Ltd, Shenzhen, China Correspondence: tfanac@cse.ust.hk
@@ 1 Introduction
The emergence of Large Language Models (LLMs) has marked a revolutionary shift in artificial intelligence, significantly transforming our understanding of natural language processing capabilities. The advent of cutting-edge LLMs like ChatGPT  OpenAI (2022), Gemma2 Team et al. (2024), and LLaMa2 Touvron et al. (2023) with their billions of parameters, has sparked the imagination of both researchers and practitioners, owing to their exceptional performance across diverse text generation tasks. Despite their widespread success in various general NLP tasks, LLMs face challenges that hinder their adoption in domain-specific applications Kang et al. (2023); Fan et al. (2023, 2024b). The primary challenges include domain-specific knowledge Privacy, constrained computing resources, and mutual knowledge transfer between the LLM and SLMs. A significant challenge arises from the inherent model heterogeneity between the LLM and SLMs, particularly when aligning distributions of output logits. The mismatch between the tokenizers of different LLM and SLMs poses a notable obstacle. Furthermore, the mutual transfer of knowledge between the server’s LLM and clients’ SLMs remains a largely unexplored area in academic literature, warranting further investigation.
To fill these gaps, we propose FedMKT, a novel federated mutual knowledge transfer framework designed to improve the performance of both large and small language models. By leveraging the complementary strengths of federated learning and knowledge distillation, FedMKT facilitates effective mutual knowledge transfer between clients’ SLMs and the LLM owned by the server.
As illustrated in Figure 1, FedMKT deploys an LLM on the server and a set of $K$ heterogeneous SLMs across various clients. The cornerstone of FedMKT lies in its selective mutual knowledge transfer process. During each round of federated learning, the clients transmit the output logits of their updated SLMs on the public dataset to the server. Subsequently, the server selectively aggregates and extracts the knowledge encoded within these SLMs output logits into the server-side LLM. This process allows the server LLM to incorporate the domain-specific knowledge learned by the clients, thereby enhancing its comprehensive capabilities. Simultaneously, the server-side LLM also selectively distills its knowledge to the clients’ SLMs, which is similar to the knowledge transfer from clients to the server. By leveraging the knowledge of the server LLM, the clients’ SLMs are able to improve their performance and generalize better to unseen data. To address the model heterogeneity between the LLM and SLMs, FedMKT incorporates a token alignment technique utilizing minimum edit distance (MinED) prior to knowledge transfer. This alignment ensures seamless integration and efficient knowledge transfer between LLM and SLMs.
@@
  • Federated Mutual Knowledge Transfer Framework. FedMKT introduces a novel federated mutual knowledge transfer framework that enables effective knowledge transfer between an LLM deployed on the server and SLMs residing on clients. This framework fills the gap by simultaneously enhancing both the server’s LLM and the clients’ SLMs. Our work is tailored for text generation tasks within the context of LLMs and supports both Heterogeneous and Homogeneous scenarios between client SLMs. To our best knowledge, our work is the first published research in this field.
@@
  • Empirical Evaluation and Performance Enhancement. Extensive experiments conducted based on various publicly available LLMs and SLMs, have shown that FedMKT exhibits competitive performance across a broad spectrum of NLP text-generation tasks. We evaluate FedMKT with Heterogeneous, Homogeneous, and One-to-One settings. The results show that the performance of SLMs can be significantly enhanced with the help of the LLM, while the LLM can deliver comparable results to fine-tuning with all clients’ data centralized.
@@ 2.1 Model Heterogeneous Federated Learning
Model heterogeneous federated learning (MHFL) aims to address the challenges associated with heterogeneity in federated learning (FL) Yang et al. (2019); McMahan et al. (2017); Liu et al. (2021); Cheng et al. (2021); Fan et al. (2024a). Initial research in MHFL primarily concentrated on addressing heterogeneity in model architectures. Several methods have been introduced to accommodate clients with different model architectures participating in a federated learning task. These methods typically involve techniques such as knowledge distillation Hinton et al. (2015), mutual learning and split learning that can handle heterogeneous models. Knowledge distillation-based MHFL methods, such as FedMD Li and Wang (2019) and FedET Cho et al. (2022), involve the server aggregating the output logits of different clients’ heterogeneous models on a public dataset to construct global logits. Mutual learning-based MHFL, such as Deep Mutual Learning (DML) Zhang et al. (2018), PFML Yang et al. (2021) and FedLoRA Yi et al. (2023), design a small homogeneous model and a large heterogeneous model in each client. Split learning-based MHFL approaches, such as FedClassAvg Jang et al. (2022) and CHFL Liu et al. (2022), share a homogeneous classifier to improve model classification while personalizing the local feature extractor.
@@ 2.2 Federated Learning for LLMs
Parameter-Efficient Fine-Tuning (PEFT) methods Houlsby et al. (2019); He et al. (2021); Lester et al. (2021); Li and Liang (2021); Hu et al. (2021) offer a direct solution to the issues of communication overhead and fine-tuning costs in federated learning (FL) for LLMs. A number of studies have built upon PEFT methods in the context of FL for LLMs, including FedPETuning Zhang et al. (2022b), Federated Adapter Tuning Cai et al. (2022), Federated Prompt Tuning Zhao et al. (2022), and FATE-LLM  Fan et al. (2023, 2024c). These findings indicate that FL clients, especially those with constrained computing and storage resources such as certain devices, can significantly profit from from PEFT approaches. These techniques facilitate the sharing of LLMs across diverse tasks, while necessitating the storage and updating of only a small number of parameters for each task, thereby reducing the overall computational and storage requirements. By leveraging PEFT methods, FL clients can efficiently adapt LLMs to their specific needs while minimizing communication overhead and fine-tuning costs.
@@ 3 The Proposed FedMKT Method
In this section, we introduce FedMKT, an innovative and parameter-efficient federated mutual knowledge transfer approach for large and small language models. The FedMKT primarily comprises two key modules: Bidirectional Token Alignment and Selective Mutual Knowledge Transfer. We will elaborate on these two key modules in Section 3.2 and Section 3.3, respectively after we define the problem we try to address in Section 3.1.
@@ 3.1 Problem Definition
We consider the federated learning setting, involving one server that owns an LLM $f_{\psi}$ parameterized by $\psi$ and $K$ clients that each client $k$ has an SLM $g_{\phi_{k}}$ parameterized by ${\phi}_{k}$. Each client owns a local private dataset denoted as $\mathcal{D}_{k}$ containing $N$ training samples, and all clients and server have access to a shared public dataset $\mathcal{D}_{p}$.
@@
FedMKT solves the optimization problems formulated in Eq.(1), Eq.(2), and Eq.(3) in an efficient and privacy-preserving manner. We illustrate the workflow of FedMKT in Figure 1 and Appendix A, elaborate on the associated training algorithm in Algorithm 1.
@@ 3.2 Bidirectional Token Alignment
To tackle this issue, we adopt dynamic programming techniques to promote robust alignment, as evidenced in studies Wan et al. (2024); Fu et al. (2023). Utilizing LLaMa2 and Bloom as illustrative examples, we establish an optimized vocabulary mapping table based on minimum edit distance(MinED). This mapping table identifies the closest Bloom token for each LLaMa2 token (e.g., ’utilize’ for ’util’). We then tokenize a sentence using both tokenizers and apply a dynamic programming algorithm to determine the optimal matching path. When multiple LLaMa2 tokens align to a single Bloom token (e.g., ’util’ and ’ize’ aligning to ’utilize’), we handle them according to the mapping table. Please see Appendix B for further details.
@@ 3.3 Selective Mutual Knowledge Transfer Between LLM and SLMs
where $\ell_{\text{CE}}$ represents the cross-entropy loss, and the model parameters $\psi$ and $\phi_{k}$ are frozen during fine-tuning.
Next, the server and clients conduct selective knowledge transfer to each other. The motivation for applying selective knowledge transfer is that some clients’ knowledge may adversely affect the performance of LLM on the server and vice versa in a heterogeneous environment. Therefore, it is critical to guarantee that the knowledge exchanged between the server and clients is positive to the performance of LLM and SLMs. To achieve this goal, we propose a selective knowledge transfer strategy on both the server and client sides, termed DualMinCE.
@@
where $\lambda$ is the hyperparameter that regulates the significance of mutual knowledge transfer.
@@ 4.1 Setup
Datasets. We evaluate FedMKT on 6 QA datasets and 2 instruction-following datasets. Specifically, for QA tasks, we use RTE Wang et al. (2019), WTC Wang et al. (2019), BoolQ Clark et al. (2019), CQA  Talmor et al. (2018), ARC-E and ARC-C Clark et al. (2018) to evaluate FedMKT. As for instruction-following tasks, we evaluate FedMKT on S-NI Wang et al. (2022), DialogSum Chen et al. (2021).
@@
  • Centralized, in which the server’s LLM is fine-tuned locally using the datasets combining private datasets of involved clients and the public dataset. In the One-to-One setting, the data of one client and the public data are used to fine-tune the server’s LLM, whereas in other settings, the data of all four clients and the public data are utilized to fine-tune the LLM;
@@
  • Standalone, where each client fine-tunes its own SLM independently using its private dataset;
  • FedAvgMcMahan et al. (2017), representing the standard federated averaging algorithm. FedAvg is only used in homogeneous settings because it requires all clients’ models have the same architecture.
@@
Evaluation Metrics. For the QA datasets, we primarily use Accuracy as the metric for evaluation, whereas for the instruction-following datasets, we primarily rely on Rouge-L. It’s worth noting that in our experiments, all methods across the three scenarios undergo zero-shot evaluation, and we utilize the lm-evaluation-harness packageGao et al. (2023) for evaluation purposes.
@@ 4.2 Evaluation on Heterogeneous Setting
Tables 2 show that FedMKT performs superior over Zero-Shot and Standalone on all clients’ SLMs. Take the RTE dataset as an example, FedMKT outperforms Zero-Shot by 34% and Standalone by 7% in relative terms on the GPT-2-xlarge SLM; FedMKT surpasses Zero-Shot by 25% and Standalone by 5% on the OPT-1.3B SLM; FedMKT-SLM achieves a 17% improvement over Zero-Shot and a 6% improvement over Standalone on the Bloom-1.1B SLM; FedMKT-SLM outperforms Zero-Shot by 18% and Standalone by 6% on the LLaMa2-1.3B SLM. These empirical results demonstrate that, by leveraging FedMKT, SLMs are able to effectively leverage the knowledge transferred from the LLM, leading to enhanced model capabilities.
@@ 4.3 Evaluation on Homogeneous Setting
The bottom sub-table of Table 3 shows that FedMTK performs better than the Zero-Shot, Standalone, and FedAvg due to the assistance of the server’s LLM. For example, in the CQA dataset, FedMKT outperforms FedAvg by 4% in relative terms on the LLaMa2-1.3 SLM and by 5% on the OPT-1.3B SLM, respectively.
@@ 5 Conclusions
In this study, we have presented FedMKT, a parameter-efficient federated mutual knowledge transfer framework tailored for LLMs and SLMs. FedMKT bridges the gap between the server-side LLMs and clients’ SLMs, enabling selective mutual knowledge transfer while preserving data privacy. Through extensive experiments across three distinct scenarios, we have demonstrated that FedMKT simultaneously boosts the performance of both LLMs and SLMs.
@@
In this study, we transfer knowledge between the server and clients using logits of a public dataset, motivated by efficiency and privacy considerations. Although empirical evidence suggests that sharing logits of public datasets between the server and clients is more privacy-preserving than sharing model gradients or parameters Li and Wang (2019); Cho et al. (2022), there is no theoretical guarantee that this approach does not compromise the privacy of clients’ sensitive data. This issue warrants further investigation. Additionally, the study does not address the potential presence of spammer Zhu et al. (2012) among clients, which could negatively impact the performance of LLM in the server and the SLMs of other clients. Detecting spammers in the FedMKT setting is identified as an important research direction. Furthermore, our study is limited by computational and storage constraints, which preclude the exploration of larger language models. This highlights the inherent trade-off between utility and efficiency. Our future research aims to investigate and optimize this trade-off.